# Directional Sheaf Hypergraph Networks: Unifying Learning on Directed and Undirected Hypergraphs

### v1

@@ 3.2 Directed Sheaf Hypergraph Laplacian
We can expand Eq. 3 by considering the special case where $q=\frac{1}{4}$. In this case, each entry of ${L}^{\vec{\mathcal{F}}}$ can be expressed to explicitly highlight the impact of the directionality of each hyperedge:
$(L^{\vec{\mathcal{F}}})_{uv}=\left\{\begin{array}[]{ll}\displaystyle\sum_{e:u\in e}\!\left(1-\frac{1}{\delta_{e}}\right)\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e},&u=v,\\ \displaystyle-\!\!\!\!\!\!\!\sum_{\begin{subarray}{c}e\in E\\ u,v\in H(e)\\ \lor u,v\in T(e)\end{subarray}}\!\!\!\!\!\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}-\mathbf{i}\left(\sum_{\begin{subarray}{c}e\in E\\ u\in T(e)\\ \land v\in H(e)\end{subarray}}\!\!\!\!\!\!\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}-\!\!\!\!\!\!\sum_{\begin{subarray}{c}e\in E\\ u\in H(e)\\ \land v\in T(e)\end{subarray}}\!\!\!\!\!\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}\right),&u\neq v.\end{array}\right$. (7)
The entry $(L^{\vec{\mathcal{F}}})_{uv}$ is determined by all hyperedges $e\in E$ that contain both nodes $u$ and $v$. From the second case in Eq. (7), whenever both $u$ and $v$ are both heads ($u,v\in H(e)$) or tails ($u,v\in T(e)$), the contribution to the real part $\Re\left((L^{\vec{\mathcal{F}}})_{uv}\right)$ is negative and given by the opposite of the normalized weight $\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}$. In the undirected case, this is the only possible contribution, which matches the expected behavior of undirected hypergraph Sheaf Laplacians, where $u,v\in T(e)$ (or, equivalently, $u,v\in H(e)$). Hyperedges where $u,v$ take opposite roles contribute to the imaginary part with their weight either negatively (if $u\in T(e)$ and $v\in H(e)$) or positively (if $u\in H(e)$ and $v\in T(e)$). Due to this, in the special case where $q=\frac{1}{4}$, $\Im\left((L^{\vec{\mathcal{F}}})_{uv}\right)$ depends on the net contribution of $u$ and $v$ across all the directed hyperedges that contain them. This is in line with the “net flow” behavior observed in directed graphs (Fiorini et al., 2023) and hypergraphs (Fiorini et al., 2024).
When interpreted as a linear operator acting on a complex signal $x\in\mathbb{C}^{nd}$, $L^{\vec{\mathcal{F}}}$ reads:
$\bigl(L^{\vec{\mathcal{F}}}(x)\bigr)_{u}=\sum_{e\in E:\,u\in e}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\sum_{\begin{subarray}{c}v\in e:\\ v\neq u\end{subarray}}\left(\vec{\mathcal{F}}_{u\trianglelefteq e}\,x_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,x_{v}\right).$ (8)
We define the Normalized Directed Sheaf Hypergraph Laplacian $L^{\vec{\mathcal{F}}}_{N}$ as:
$L^{\vec{\mathcal{F}}}_{N}:=D_{V}^{-\frac{1}{2}}L^{\vec{\mathcal{F}}}D_{V}^{-\frac{1}{2}}.$
@@
$L^{\vec{\mathcal{F}}}_{N}=D_{V}^{-\frac{1}{2}}\underbrace{(D_{V}-Q^{\vec{\mathcal{F}}})}_{L^{\vec{\mathcal{F}}}}D_{V}^{-\frac{1}{2}}=I_{nd}-Q^{\vec{\mathcal{F}}}_{N},\;\;\text{ where }\;\;Q^{\vec{\mathcal{F}}}_{N}:=D_{V}^{-\frac{1}{2}}{B^{(q)}}^{\dagger}D_{E}^{-1}B^{(q)}D_{V}^{-\frac{1}{2}}.$ (9)
@@ 3.3 Spectral Properties
In this section, we prove that our proposed Normalized Directed Sheaf Hypergraph Laplacian $L^{\vec{\mathcal{F}}}_{N}$ enjoys all the spectral properties that are required of a principled convolutional operator. Specifically, we show that $L^{\vec{\mathcal{F}}}_{N}$ is diagonalizable, has real eigenvalues, is positive semidefinite, and has a bounded spectrum. The proofs of the claims in this and the next section are provided in Appendix A.
We begin by showing that $L^{\vec{\mathcal{F}}}_{N}$ admits an eigenvalue decomposition with real eigenvalues: {restatable}theoremdiagonalizable $L^{\vec{\mathcal{F}}}_{N}$ is diagonalizable with real eigenvalues.
Next, we derive the formula of the Dirichlet energy function associated with $L^{\vec{\mathcal{F}}}_{N}$, which provides a measure of the global smoothness of a signal $x\in\mathbb{C}^{nd}$ across the entire hypergraph: {restatable}theoremdirichlet The Dirichlet energy induced by $L^{\vec{\mathcal{F}}}_{N}$ for a signal $x\in\mathbb{C}^{nd}$ is:
$\mathcal{E}_{N}(x)=x^{\dagger}L^{\vec{\mathcal{F}}}_{N}x=\frac{1}{2}\sum_{e\in E}\frac{1}{\delta_{e}}\sum_{\begin{subarray}{c}u,v\in e:\\ u\neq v\end{subarray}}\Bigl\|\vec{\mathcal{F}}_{u\trianglelefteq e}\,D_{u}^{-\frac{1}{2}}x_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,D_{v}^{-\frac{1}{2}}x_{v}\Bigr\|_{2}^{2}.$
By leveraging the two previous theorems, we show that the spectrum of the Directed Sheaf Hypergraph Laplacian $L^{\vec{\mathcal{F}}}_{N}$ only admits (real) non-negative eigenvalues: {restatable}corollarypsd $L^{\vec{\mathcal{F}}}_{N}$ is positive semidefinite.
Finally, we prove that the spectrum of $L^{\vec{\mathcal{F}}}_{N}$ is upper bounded by 1: {restatable}theorembounded $\lambda_{\max}(L^{\vec{\mathcal{F}}}_{N})\leq 1$.
@@ 3.4 Generalization Properties
First, we discuss the relationship between $L^{\vec{\mathcal{F}}}$ and the Sheaf Laplacian introduced by Hansen and Gebhart (2020) and highlight its connection with the classical graph Laplacian defined as L:=DAL:=D-A where $A$ is the graph’s adjacency matrix and $D$ its degree matrix (see (Biggs, 1993) for a reference) for the undirected case: {restatable}theorembodnar For a $2$-uniform hypergraph without directions, the Laplacian operator $L^{\vec{\mathcal{F}}}$ reduces to the Sheaf Laplacian (Hansen and Gebhart, 2020) (up to a scaling factor of 2) and, when considering the case of a trivial Sheaf (where $\mathcal{F}_{u\trianglelefteq e}=1$), it coincides with the classical graph Laplacian (up to a scaling factor of 2).
@@
With Section 3.4, we show that $L^{\vec{\mathcal{F}}}$ generalizes several Laplacians designed for directed graphs like the Magnetic Laplacian (Zhang et al., 2021) and the Sign-Magnetic Laplacian (Fiorini et al., 2023): {restatable}theoremmagnet For a directed $2$-uniform hypergraph with unitary edge weights (i.e., $w_{e}=1,e\in E$) containing both directed and undirected edges, $L^{\vec{\mathcal{F}}}$ recovers, as a special case, the Magnetic Laplacian (Zhang et al., 2021) for any $q\in\mathbb{R}$ and the Sign-Magnetic Laplacian (Fiorini et al., 2023) when $q=\frac{1}{4}$.
@@
In the context of hypergraphs, our operator naturally recovers existing hypergraph Laplacians. We begin by showing that it recovers the undirected hypergraph Laplacian of Zhou et al. (2006): {restatable}theoremzhou Given a hypergraph $\mathcal{H}$ (directed or undirected), the normalized Directed Hypergraph Laplacian $L^{\vec{\mathcal{F}}}_{N}$ recovers, as a special case, the undirected hypergraph Laplacian of Zhou et al. (2006). Lastly, we show that $L^{\vec{\mathcal{F}}}$ generalizes the Generalized Directed Laplacian introduced by Fiorini et al. (2024): {restatable}theoremgedi Given a directed hypergraph $\mathcal{H}$ with unitary weights associated to each hyperedge (i.e., $w_{e}=1$), the Normalized Directed Sheaf Hypergraph Laplacian $L^{\vec{\mathcal{F}}}_{N}$ recovers, as a special case, the Generalized Directed Laplacian $\vec{L}_{N}$ of Fiorini et al. (2024).
Crucially, our Laplacian operator does not recover the (linear) Sheaf Hypergraph Laplacian proposed by Duta et al. (2023). This is intentional, since Duta et al. (2023)’s formulation fails to satisfy the spectral properties that are required to build a well-defined convolutional operator, most notably positive semidefiniteness (which only holds in the the 2-uniform case). This is better described in Appendix D. Consequently, when either setting $q=0$ or considering undirected hypergraphs, our operator enjoys all the properties that are required of a principled Laplacian matrix and, thus, it provides (to the best of our knowledge) the first definition of a Sheaf Hypergraph Laplacian suitable for undirected hypergraphs.
@@ 4 Directional Sheaf Hypergraph Network
In this section, we describe our proposed Directional Sheaf Hypergraph Network (DSHN). Inspired by Hansen and Gebhart (2020), we define the sheaf diffusion process on a directed hypergraph $\mathcal{H}$ as an extension of classical heat diffusion, which plays a central role in spectral-bases GNN convolution operators (Kipf and Welling, 2017). Starting from the differential equation:
$\dot{X}_{t}=-\,L^{\vec{\mathcal{F}}}_{N}X_{t}.$
@@
$X_{t+1}=X_{t}-L^{\vec{\mathcal{F}}}_{N}X_{t}=(I_{nd}-L^{\vec{\mathcal{F}}}_{N})X_{t}.$
@@
$X_{t+1}=\sigma\left((I_{nd}-L^{\vec{\mathcal{F}}}_{N})\,(I_{n}\otimes W_{1})\,X_{t}\,W_{2}\right)=\sigma\left(Q^{\vec{\mathcal{F}}}_{N}\,(I_{n}\otimes W_{1})\,X_{t}\,W_{2}\right)\in\mathbb{C}^{nd\times f},$ (10)
where $W_{1}\in\mathbb{R}^{d\times d}$ and $W_{2}\in\mathbb{R}^{f\times f}$ are trainable weight matrices, and $\otimes$ denotes the Kronecker product. $X_{0}$ is obtained from the input matrix of node features of size $n\times f$, to which a linear projection is applied to produce an $n\times(df)$ matrix, which is then reshaped into a $(nd)\times f$ matrix before applying the diffusion process. Finally, since the convolutional layer operates in the complex domain, we transform the output of the final convolutional layer into a real-valued representation right before feeding it to the classification head. Following Zhang et al. (2021); Fiorini et al. (2023), we do so by applying an unwind operation, concatenating the real and imaginary components of the complex features as follows:
$\mathrm{unwind}(X)=\Re(X)\,\|\,\Im(X)\in\mathbb{R}^{n\times 2f}.$
@@
The expressive power of sheaf-theoretical approaches lies in their ability to define a diffusion operator via a learnable $d\times d$ restriction map associated to each node–edge pair. Following Bodnar et al. (2022), we learn each restriction map as a function of the corresponding node and hyperedge features. In particular, for each hyperedge $e\in E$ and each node $u\in e$, each directionless restriction map $\mathcal{F}_{v\trianglelefteq e}$ is parametrized as $\mathcal{F}_{v\trianglelefteq e}=\Phi(x_{u}\,\|\,x_{e})\in\mathbb{R}^{d\times d}$, where $x_{u}$ is the node feature and $x_{e}$ is the hyperedge feature and $\Phi$ is an MLP. If hyperedge features are not explicitly provided, they are computed by aggregating the features of the hyperedge’s nodes via a mean or a sum. Given that the node and the hyperedge features $x_{v}$ and $x_{e}$ are complex-valued due to Eq. 10, we employ the unwind operation to map them into a form suitable for input to $\Phi$. Activation functions can be either sigmoid or tanh.
@@
A key objective in our model design is to balance predictive performance with computational efficiency. Constructing an $nd\times nd$ Laplacian matrix from $d\times d$ restriction maps inherently increases complexity, an issue already noted in the graph setting (Bodnar et al., 2022). To mitigate this, we introduce DSHNLight, a variant of DSHN that achieves competitive, and in some cases superior, results across several datasets (Tables 1 and 3) at a significantly lower computational cost (see Section E.1). In it, we detach the gradient computation during the Laplacian’s construction: this way, the model continues to rely on the predicted restriction maps, but avoids costly gradient propagation. In DSHNLight, the parameters of the MLP responsible for predicting the restriction maps (which are encoded in $\Phi(\cdot)$) remain fixed throughout the training process. The model’s adaptability arises from the initial projection layer, which embeds the inputs into a shared feature space where they can be more effectively processed through these parameters (see Appendix E for a better visualization). This phenomenon aligns with insights from the literature on overparameterization (Jacot et al., 2018; Arora et al., 2019) and extreme learning machines (Huang et al., 2006), where fixed random projections can still yield strong generalization due to the expressive power of the input embeddings. Further details on the difference between the two approaches and their asymptotic complexity at inference time are discussed in Section E.1

### v2

@@ 3.2 Directed Sheaf Hypergraph Laplacian
We can expand Eq. 3 by considering the special case where $q=\frac{1}{4}$. In this case, each entry of $\textbf{L}^{\vec{\mathcal{F}}}$ can be expressed to explicitly highlight the impact of the directionality of each hyperedge:
$(\mathbf{L}^{\vec{\mathcal{F}}})_{uv}=\left\{\begin{array}[]{ll}\displaystyle\sum_{e:u\in e}\!\left(1-\frac{1}{\delta_{e}}\right)\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e},&u=v,\\ \displaystyle-\!\!\!\!\!\!\!\sum_{\begin{subarray}{c}e\in E\\ u,v\in H(e)\\ \lor u,v\in T(e)\end{subarray}}\!\!\!\!\!\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}-i\left(\sum_{\begin{subarray}{c}e\in E\\ u\in T(e)\\ \land v\in H(e)\end{subarray}}\!\!\!\!\!\!\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}-\!\!\!\!\!\!\sum_{\begin{subarray}{c}e\in E\\ u\in H(e)\\ \land v\in T(e)\end{subarray}}\!\!\!\!\!\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}\right),&u\neq v.\end{array}\right$. (7)
The entry $(\mathbf{L}^{\vec{\mathcal{F}}})_{uv}$ is determined by all hyperedges $e\in E$ that contain both nodes $u$ and $v$. From the second case in Eq. (7), whenever both $u$ and $v$ are both heads ($u,v\in H(e)$) or tails ($u,v\in T(e)$), the contribution to the real part $\Re\left((\mathbf{L}^{\vec{\mathcal{F}}})_{uv}\right)$ is negative and given by the opposite of the normalized weight $\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}$. In the undirected case, this is the only possible contribution, which matches the expected behavior of undirected hypergraph Sheaf Laplacians, where $u,v\in T(e)$ (or, equivalently, $u,v\in H(e)$). Hyperedges where $u,v$ take opposite roles contribute to the imaginary part with their weight either negatively (if $u\in T(e)$ and $v\in H(e)$) or positively (if $u\in H(e)$ and $v\in T(e)$). Due to this, in the special case where $q=\frac{1}{4}$, $\Im\left((\mathbf{L}^{\vec{\mathcal{F}}})_{uv}\right)$ depends on the net contribution of $u$ and $v$ across all the directed hyperedges that contain them. This is in line with the “net flow” behavior observed in directed graphs (Fiorini et al., 2023) and hypergraphs (Fiorini et al., 2024).
When interpreted as a linear operator acting on a complex signal $\mathbf{x}\in\mathbb{C}^{nd}$, $\mathbf{L}^{\vec{\mathcal{F}}}$ reads:
$\bigl(\mathbf{L}^{\vec{\mathcal{F}}}(\mathbf{x})\bigr)_{u}=\sum_{e\in E:\,u\in e}\frac{1}{\delta_{e}}\,\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\sum_{\begin{subarray}{c}v\in e:\\ v\neq u\end{subarray}}\left(\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{x}_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,\mathbf{x}_{v}\right).$ (8)
We define the Normalized Directed Sheaf Hypergraph Laplacian $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ as:
$\mathbf{L}^{\vec{\mathcal{F}}}_{N}:=\mathbf{D}_{V}^{-\frac{1}{2}}\mathbf{L}^{\vec{\mathcal{F}}}\mathbf{D}_{V}^{-\frac{1}{2}}.$
@@
$\mathbf{L}^{\vec{\mathcal{F}}}_{N}=\mathbf{D}_{V}^{-\frac{1}{2}}\underbrace{(\mathbf{D}_{V}-\mathbf{Q}^{\vec{\mathcal{F}}})}_{\mathbf{L}^{\vec{\mathcal{F}}}}\mathbf{D}_{V}^{-\frac{1}{2}}=\textbf{I}_{nd}-\mathbf{Q}^{\vec{\mathcal{F}}}_{N},\;\;\text{ where }\;\;\mathbf{Q}^{\vec{\mathcal{F}}}_{N}:=\mathbf{D}_{V}^{-\frac{1}{2}}{\mathbf{B}^{(q)}}^{\dagger}\mathbf{D}_{E}^{-1}\mathbf{B}^{(q)}\mathbf{D}_{V}^{-\frac{1}{2}}.$ (9)
@@ 3.3 Spectral Properties
We now establish that our proposed Normalized Directed Sheaf Hypergraph Laplacian $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ satisfies all the spectral properties required for a principled convolutional operator. Specifically, we show that $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ is diagonalizable, has real, nonnegative eigenvalues, is positive semidefinite, and has a bounded spectrum. These ensure that the Fourier transform is well-defined and that polynomial filters of $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ implement localized, stable convolutions, in direct analogy with classical spectral-based approaches (Shuman et al., 2013; Kipf and Welling, 2017; Defferrard et al., 2016b). Failure to preserve them can result in a complete breakdown of the connection between message passing, the Fourier transform of a (hyper-)graph and its connection with signal theory. The proofs of the claims in this and the next section are provided in Appendix A.
We begin by showing that $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ admits an eigenvalue decomposition with real eigenvalues: {restatable}theoremdiagonalizable $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ is diagonalizable with real eigenvalues.
Next, we derive the formula of the Dirichlet energy function associated with $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$, which provides a measure of the global smoothness of a signal $x\in\mathbb{C}^{nd}$ across the entire hypergraph: {restatable}theoremdirichlet The Dirichlet energy induced by $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ for a signal $\mathbf{x}\in\mathbb{C}^{nd}$ is:
$\mathcal{E}_{N}(\mathbf{x})=\mathbf{x}^{\dagger}\mathbf{L}^{\vec{\mathcal{F}}}_{N}\mathbf{x}=\frac{1}{2}\sum_{e\in E}\frac{1}{\delta_{e}}\sum_{\begin{subarray}{c}u,v\in e:\\ u\neq v\end{subarray}}\Bigl\|\vec{\mathcal{F}}_{u\trianglelefteq e}\,\mathbf{D}_{u}^{-\frac{1}{2}}\mathbf{x}_{u}-\vec{\mathcal{F}}_{v\trianglelefteq e}\,\mathbf{D}_{v}^{-\frac{1}{2}}\mathbf{x}_{v}\Bigr\|_{2}^{2}.$
By leveraging the two previous theorems, we show that the spectrum of the Directed Sheaf Hypergraph Laplacian $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ only admits (real) non-negative eigenvalues: {restatable}corollarypsd $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ is positive semidefinite.
Finally, we prove that the spectrum of $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ is upper bounded by 1: {restatable}theorembounded $\lambda_{\max}(\mathbf{L}^{\vec{\mathcal{F}}}_{N})\leq 1$.
@@ 3.4 Generalization Properties
First, we discuss the relationship between $\mathbf{L}^{\vec{\mathcal{F}}}$ and the Sheaf Laplacian introduced by Hansen and Gebhart (2020) and highlight its connection with the classical graph Laplacian defined as $\textbf{L}:=\textbf{D}-\textbf{A}$ where $A$ is the graph’s adjacency matrix and $D$ its degree matrix (see (Biggs, 1993) for a reference) for the undirected case: {restatable}theorembodnar For a $2$-uniform hypergraph without directions, the Laplacian operator $\mathbf{L}^{\vec{\mathcal{F}}}$ reduces to the Sheaf Laplacian (Hansen and Gebhart, 2020) (up to a scaling factor of 2) and, when considering the case of a trivial Sheaf (where $\mathcal{F}_{u\trianglelefteq e}=1$), it coincides with the classical graph Laplacian (up to a scaling factor of 2).
@@
With Section 3.4, we show that $\mathbf{L}^{\vec{\mathcal{F}}}$ generalizes several Laplacians designed for directed graphs like the Magnetic Laplacian (Zhang et al., 2021) and the Sign-Magnetic Laplacian (Fiorini et al., 2023): {restatable}theoremmagnet For a directed $2$-uniform hypergraph with unitary edge weights (i.e., $w_{e}=1,e\in E$) with directed and undirected edges, $\mathbf{L}^{\vec{\mathcal{F}}}$ recovers, as a special case, the Magnetic Laplacian (Zhang et al., 2021) for any $q\in\mathbb{R}$ and the Sign-Magnetic Laplacian (Fiorini et al., 2023) when $q=\frac{1}{4}$.
@@
In the context of hypergraphs, our operator naturally recovers existing hypergraph Laplacians. We begin by showing that it recovers the undirected hypergraph Laplacian of Zhou et al. (2006): {restatable}theoremzhou Given a hypergraph $\mathcal{H}$ (directed or undirected), the normalized Directed Hypergraph Laplacian $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ recovers, as a special case, the undirected hypergraph Laplacian of Zhou et al. (2006). We show that $\mathbf{L}^{\vec{\mathcal{F}}}$ generalizes the Generalized Directed Laplacian proposed by Fiorini et al. (2024): {restatable}theoremgedi Given a directed hypergraph $\mathcal{H}$ with unitary weights associated to each hyperedge (i.e., $w_{e}=1$), the Normalized Directed Sheaf Hypergraph Laplacian $\mathbf{L}^{\vec{\mathcal{F}}}_{N}$ recovers, as a special case, the Generalized Directed Laplacian $\vec{\mathbf{L}}_{N}$ of Fiorini et al. (2024).
Crucially, our Laplacian operator does not recover the (linear) Sheaf Hypergraph Laplacian proposed in (Duta et al., 2023). This difference is intentional: the formulation proposed in their work, defined for any pair of nodes $u,v\in V$ as:
$(\mathbf{L}^{\mathcal{F}})_{uu}=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e},\qquad(\mathbf{L}^{\mathcal{F}})_{uv}=-\!\sum_{\begin{subarray}{c}e:u,v\in e\\ v\neq u\end{subarray}}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e},$ (10)
fails to satisfy several key spectral properties required to construct a well-defined convolutional operator, most notably positive semidefiniteness. Comparing this definition with Eq. 3, it is evident that, for undirected hypergraphs or when $q=0$, the two operators differ in the diagonal term: our formulation introduces the factor $(1-\tfrac{1}{\delta_{e}})$ rather than $\tfrac{1}{\delta_{e}}$. Their definition suffices to guarantee such properties only in the special case of 2-uniform hypergraphs (i.e, standard graphs), however this discrepancy can produce negative eigenvalues for general cases, preventing the matrix from being interpreted as a diffusion operator and precluding the definition of a stable Fourier transform. Polynomial filters based on the Laplacian of Duta et al. (2023) therefore cannot guarantee localized and stable convolutions, limiting its general applicability. A more detailed analysis, including an example that demonstrates this drawback, is presented in Appendix E.
In contrast, when setting $q=0$ or considering undirected hypergraphs, our operator satisfies all the properties required of a principled Laplacian matrix and, to the best of our knowledge, provides the first definition of a Sheaf Hypergraph Laplacian suitable for undirected hypergraphs.
@@ 4 Directional Sheaf Hypergraph Network
In this section, we describe our proposed Directional Sheaf Hypergraph Network (DSHN). Inspired by Hansen and Gebhart (2020), we define the sheaf diffusion process on a directed hypergraph $\mathcal{H}$ as an extension of classical heat diffusion, which plays a central role in spectral-based GNN convolution operators (Kipf and Welling, 2017). Starting from the differential equation:
$\dot{\mathbf{X}}_{t}=-\,\mathbf{L}^{\vec{\mathcal{F}}}_{N}\mathbf{X_{t}}.$
@@
$\mathbf{X_{t+1}}=\mathbf{X_{t}}-\mathbf{L}^{\vec{\mathcal{F}}}_{N}\mathbf{X_{t}}=(\mathbf{I}_{nd}-\mathbf{L}^{\vec{\mathcal{F}}}_{N})\mathbf{X_{t}}.$
@@
$\mathbf{X}_{t+1}=\sigma\left((\mathbf{I}_{nd}-\mathbf{L}^{\vec{\mathcal{F}}}_{N})\,(\mathbf{I}_{n}\otimes\mathbf{W_{1}})\,\mathbf{X}_{t}\,\mathbf{W}_{2}\right)=\sigma\left(\mathbf{Q}^{\vec{\mathcal{F}}}_{N}\,(\mathbf{I}_{n}\otimes\mathbf{W}_{1})\,\mathbf{X}_{t}\,\mathbf{W}_{2}\right)\in\mathbb{C}^{nd\times f},$ (11)
where $\mathbf{W_{1}}\in\mathbb{R}^{d\times d}$ and $\mathbf{W_{2}}\in\mathbb{R}^{f\times f}$ are trainable weight matrices, and $\otimes$ denotes the Kronecker product. $\mathbf{X}_{0}$ is obtained from the input matrix of node features of size $n\times f$, to which a linear projection is applied to produce an $n\times(df)$ matrix, which is then reshaped into a $(nd)\times f$ matrix before applying the diffusion process. Finally, since the convolutional layer operates in the complex domain, we transform the output of the final convolutional layer into a real-valued representation right before feeding it to the classification head. Following Zhang et al. (2021); Fiorini et al. (2023), we do so by applying an unwind operation, concatenating the real and imaginary components of the complex features as follows:
$\mathrm{unwind}(\mathbf{X})=\Re(\mathbf{X})\,\|\,\Im(\mathbf{X})\in\mathbb{R}^{n\times 2f}$
@@
The expressive power of sheaf-theoretical approaches lies in their ability to define a diffusion operator via a learnable $d\times d$ restriction map associated to each node–edge pair. Following Bodnar et al. (2022), we learn each restriction map as a function of the corresponding node and hyperedge features. In particular, for each hyperedge $e\in E$ and each node $u\in e$, each directionless restriction map $\mathcal{F}_{v\trianglelefteq e}$ is parametrized as $\mathcal{F}_{v\trianglelefteq e}=\Phi(\mathbf{x}_{v}\,\|\,\mathbf{x}_{e})\in\mathbb{R}^{d\times d}$, where $x_{u}$ is the node feature and $\mathbf{x}_{e}$ is the hyperedge feature and $\Phi$ is an MLP. If hyperedge features are not explicitly provided, they are computed by aggregating the features of the hyperedge’s nodes via a mean or a sum. Given that the node and the hyperedge features $\mathbf{x}_{v}$ and $\mathbf{x}_{e}$ are complex-valued due to Eq. 11, we employ the same unwind operation to map them into a form suitable for input to $\Phi$. Activation functions can be either sigmoid or tanh.
@@
A key objective in our model design is to balance predictive performance with computational efficiency. Constructing an $nd\times nd$ Laplacian matrix from $d\times d$ restriction maps inherently increases complexity, an issue already noted in the graph setting (Bodnar et al., 2022). To mitigate this, we introduce DSHNLight, a variant of DSHN that achieves competitive, and in some cases superior, results across several datasets (Tables 1 and 3) at a significantly lower computational cost (see Section C.1). In it, we detach the gradient computation during the Laplacian’s construction: this way, the model continues to rely on the predicted restriction maps, but avoids costly gradient propagation. In DSHNLight, the parameters of the MLP responsible for predicting the restriction maps (which are encoded in $\Phi(\cdot)$) remain fixed throughout the training process. The model’s adaptability arises from the initial projection layer, which embeds the inputs into a shared feature space where they can be more effectively processed through these parameters (see Appendix C for a better visualization). This phenomenon aligns with insights from the literature on overparameterization (Arora et al., 2019) and extreme learning machines (Huang et al., 2006), where fixed random projections can still yield strong generalization due to the expressive power of the input embeddings. Further details on the difference between the two approaches are discussed in Appendix C.
Computational Complexity
We provide an estimate of the asymptotic complexity of our model at inference time. Let $n$ denote the number of nodes, $m$ the number of hyperedges, $d$ the stalk dimension, $c$ the product of input and output feature dimensions in the linear transformation, $\bar{e}$ the average hyperedge size, and $\bar{v}$ the average number of hyperedges a node participates in. Summing the contributions from the feature transformation, message passing, learning of restriction maps, and Laplacian assembly, the overall complexity is $\mathcal{O}(n(c^{2}+d)+m(\bar{e}d+\bar{e}^{2}(d+c)+\bar{v}c))$ for diagonal maps, and $\mathcal{O}(n(c^{2}+d^{3})+m(\bar{e}d^{3}+\bar{e}^{2}(d^{3}+dc)+\bar{v}d^{2}c))$ for non-diagonal maps, sharing the same asymptotic complexity as SheafHyperGNN by Duta et al. (2023). For a comprehensive analysis of the contributions leading to this asymptotic complexity, we refer the reader to Section C.1.